14. A drive shaft of radius 15 cm has an angular frequency of 100 rpm. Determine

a) The tangential velocity

b) The centripetal acceleration



1
Expert's answer
2021-01-10T18:27:16-0500

Let "R = 15cm = 0.15m" be the radius, and "f = 100rpm = 5\/3Hz" be the angular frequency. By defintion, the period of motion (time of one revolution) is:


"T = \dfrac{1}{f}"

The tangential velocity is then the full circule length divided by period:


"v = \dfrac{2\pi R}{T} = 2\pi R f\\
v = 2\pi \cdot 0.15\cdot \dfrac53 \approx 1.57 m/s"

The centripetal acceleration is:


"a = \dfrac{v^2}{R}\\
a = \dfrac{1.57^2}{0.15} \approx 16.4m/s^2"

Answer. a) 1.57 m/s, b) 16.4 m/s^2.